1. The number x is multiplied by itself and the result y is less than the number x.

For example, x*x = x2 = y and y < x. 

i.   State a set name in which numbers behave in such an assertion above.

ii.     State any number in the stated set in (a) above and verify the assertion. 


1
Expert's answer
2021-03-09T07:50:42-0500

1. "y=x^2<x"

i.


"x^2<x=>x^2-x<0=>x(x-1)<0"

"=>x\in(0, 1)"

The opened interval "(0, 1)."


ii. Let "x=\dfrac{1}{2}." Then


"y(\dfrac{1}{2})=(\dfrac{1}{2})^2=\dfrac{1}{4}<\dfrac{1}{2}=x"